OpenStudy (anonymous):

what is radical 27 + radical48

OpenStudy (anonymous):

\[7\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{3*9}+\sqrt{16*3}\] \[3\sqrt{3}+4\sqrt{3}\] \[7\sqrt{3}\]